# Learning from One Continuous Video Stream

### v1

@@ Abstract
We introduce a framework for online learning from a single continuous video stream – the way people and animals learn, without mini-batches, data augmentation or shuffling. This poses great challenges given the high correlation between consecutive video frames and there is very little prior work on it. Our framework allows us to do a first deep dive into the topic and includes a collection of streams and tasks composed from two existing video datasets, plus methodology for performance evaluation that considers both adaptation and generalization. We employ pixel-to-pixel modelling as a practical and flexible way to switch between pre-training and single-stream evaluation as well as between arbitrary tasks, without ever requiring changes to models and always using the same pixel loss. Equipped with this framework we obtained large single-stream learning gains from pre-training with a novel family of future prediction tasks, found that momentum hurts, and that the pace of weight updates matters. The combination of these insights leads to matching the performance of IID learning with batch size 1, when using the same architecture and without costly replay buffers.
@@ 1 Introduction
Figure 1: Top: We introduce a framework for studying continuous learning in a single video stream. This is a natural yet unstudied problem, different from standard independent and identically distributed (IID) learning in video where batches contain clips from random videos in a random order. Bottom: We propose to employ pixel-to-pixel models to evaluate our approach across prediction tasks (prediction of future frames, depth, segmentation). We measure both adaptation to the video stream – the model here updates its weights (learns) continuously to improve prediction – as well as generalization to out-of-stream clips – the model being adapted on the first stream is now evaluated on a different held-out stream without being allowed to adapt to it. We propose to maximize both adaptation and generalization.
@@
But what happens if we attempt to learn continuously from a single video stream (Figure 1), meaning batch size 1 and high frame rate (e.g. 25 fps)? We do not see thousands of dogs per second. Different ImageNet images of even a same dog breed are quite uncorrelated at several scales, for example capturing a wide diversity of edges, textures, etc., hence standard deep learning approaches lead to smooth optimization. This is however very different from learning from a single video stream, where things happen slowly, at much longer timescales, with great correlation or even no change between frames for periods of time; see Figure 2. Do standard deep learning tools work in this setting? This is currently unknown territory, with very few attempts touching it at all [34] – and it is the focus of this paper.
Why is single-stream learning interesting in practice? Ideally, models would be able to adapt after deployment to their environment, hence receiving a single stream of information. This is needed for embodied intelligence (robotics), but also for any type of digital assistant that we would want to be able to adapt to a specific user’s needs.
@@ 2 Related work
Online learning from a single video stream. Almost all video research considers batches of shuffled video clips for training models. An inspiring but rare exception [34] investigated broadly the challenges of self-supervised learning from continuous data streams and proposed using a minimum-redundancy “replay” buffer [28] to deal with temporal correlations and learn more robust representations. Replay buffers are generally not seen as the most interesting avenue for progress in continual learning [42] and particularly in our setting as they increase computation proportionally to the size of the batch used; it would be much more appealing to improve training efficiency while retaining batch size 1 (but we have a small experiment exploring a replay buffer too). Additionally, in [34] performance on the video streams itself was never evaluated, only on unrelated image classification tasks. Differently, our evaluation sheds light into the interrelationship between learning, adaptation and generalization in the training video streams directly, on multiple video tasks, and we also analyze more factors, such as the impact of optimization. In another work  [47], the authors leverage the correlations in the data to propose a more efficient semi-supervised object detector from streaming video. Relevant for our work are also attempts to train a ConvNet [22] or a ViT [9] from a single image [1] or a single long video [44]. Encouragingly, they showed that learning can be successful (especially for the shallower layers), but still use data shuffling, augmentation and large batches, so not really learning online. Finally, one possible concern with single-stream learning is how to fully exploit available hardware parallelism in a batch size 1 setting – a few recent paper have looked into parallelization along network depth [7, 24, 25].
Online continual learning. Zooming out, our work also relates to the online continual learning literature. In this area models learn from a stream of data, visited once. One problem that has been identified in this setting is cheating by exploiting labels in the data  [16], or by learning spurious features [46] that have limited generalisation power. However, these works again do not consider the case where the input stream is highly-temporally correlated, like in real-world videos. Instead, they typically focus on streams of independent images [4].
Continual learning. Zooming out further, the continual learning literature (e.g. EWC  [20]), frequently considers the setting where a model learns a sequence of tasks, and the goal is to learn new tasks as fast as possible without forgetting previous tasks. For example, class incremental learning (e.g. iCarl  [37]) is an instance of the problem, where a model starts by learning a subset of the classes in a dataset, and new classes are added during training. Multiple benchmarks have been proposed to measure backward transfer (information retention) and forward transfer (how fast can the model learn a new task, or how generalisable are the features learnt in early tasks), such as CLEAR [23]. However, these works do not study the problem in the presence of temporal correlations. Instead they focus on the non-stationarity of the distribution, e.g. new classes being added [37], or the distribution of the input data changes significantly (e.g. different permutations of MNIST  [20]).
@@
Representation learning to the rescue. Recent continual learning papers reported that some of the challenges of the problem are partially mitigated by using features pretrained in IID settings [27, 36, 45]. Various pretraining strategies exist: self-distillation (e.g. BYOL  [14], BRaVe [38], DINO [5]), contrastive (CLIP [35], VideoMoco [29], VITO [31]), and masked auto-encoding (e.g. VideoMAE [43], TubeViT [32], AudioVisualMAE [12], SiameseMAE [15],  [2]). We investigate this aspect in our setup with temporally-correlated data and propose new future prediction pretraining approaches that transfer well to single-stream learning.
@@ 3 The Framework
We introduce here a new overall framework for exploring single-stream video learning. It consists of the following: 1) a uniform pixel-to-pixel modelling approach that streamlines using the same model in different tasks / video streams / switching between batch pretraining and learning from a single-stream, both because no change to the model structure is ever required and because the exact same loss function can always be used. 2) we create streams and tasks by adapting 2 existing public video datasets having semantic annotations 3) an approach to evaluate the specificities of single-stream learning by computing two metrics, one related to adaptation to the training stream, and another to out-of-stream generalization.
Overview. The following is repeated forever: a (potentially pretrained) model processes $n$ frames (a time step) from an input video stream, and predicts frames for another time step. Online (in-stream) performance metrics as well as an L2 loss function are then recorded by comparing predicted and target frames. Given the loss, gradients are computed and the model weights are updated by an optimizer. The target frames can come from the input stream itself or from an associated stream (e.g. for datasets having semantic segmentation or depth). In parallel with in-stream evaluation, we periodically assess performance on a held-out stream (composed of clips from the validation set of the dataset), to assess generalization.
3.1 Uniform pixel-to-pixel modelling
@@
Replay buffer. We explored replay buffers [28], which keep a large cache of previous examples, then form batches by sampling from from it. They have the disadvantage of increasing computational cost significantly over operating with batch size 1 (increases the theoretical computational cost of learning $K$ times, where $K$ is the batch size).
@@ 3.3 Evaluation
Learning a video stream task with a continuously updated model can be seen to decompose into: 1) latching on to / discovering strong features that are general, for example learning that certain configuration of edges correspond to chair legs for semantic segmentation (generalization), which is likely to happen over long timescales. 2) by specializing to the particular scene being seen in the video, e.g. learning that the chair is the only red object in the scene and using redness to label the chair pixels (adaptation), which can happen over short timescales. Adaptation can be quite useful and lead to efficient visual mechanisms, but may not necessarily lead to generalization. Dummy models that only e.g. exploit temporal correlations in the target stream could misleadingly appear to perform very well [16]; see Table 4, Blind model.
We propose to evaluate both aspects by computing a score continuously in-stream, on the video stream task the model is learning from, to measure adaptation. Separately and periodically we compute the same score out-of-stream, on a held-out stream with the optimizer disabled (out-of-stream). We use the training set of Ego4D and ScanNet to do in-stream evaluation, and the validation sets for out-of-stream evaluation.
Cumulative scores. Comparing single-stream learning models and approaches can be difficult if done naively. On the training stream, difficulty of the task may vary over time, for example certain parts of a stream may be more unpredictable due to camera motion or other factors, leading to natural oscillations. On the generalization side we care not just about performance at a particular moment in time, but also how fast it ramps up as training on the sequential video stream progresses. With this in mind, we use a global score over the whole stream – we do this by averaging the performance over 10,000 evenly-spaced points interpolated from the steps in both the in-stream and out-of-stream setting.
@@ 4 Generalized Future Prediction
One of the most promising directions for continual learning, as identified in prior work, is representation learning. We are particularly interested in pixel-to-pixel approaches keeping consistent with the overall proposed framework where we never change neither the architecture nor the loss.
To this effect we introduce a novel family of pretraining methods that generalize future video prediction [41]. All of them follow a same pattern: given one input video clip (we consider 4 frames) the model is trained to predict a future clip from the same video, also 4 frames long. We consider three variants:
  • The easiest one, Guided future prediction, overlays a few patches from the future clip into the input video clip, hence narrowing down the range of possibilities.
@@
  • Masked future prediction is a variation of Masked Auto-Encoding where the model must predict the future clip based on a partial view of the current clip. Note this is strictly harder than vanilla future prediction since the model cannot even fully see the present clip.
These variations are illustrated in Fig. 3. Guided future prediction is related to Siamese MAE [15] and Counterfactual World Modelling [2], but simpler, requiring a single forward pass over the model. Masked future prediction is related to video MAE-type approaches [43, 11] but with a strict separation between disjoint input and output clips, whereas in traditional video MAE there is a single clip that gets uniformly corrupted as part of denoising auto-encoding.
@@ 5.2 Pretraining
Figure 8: Linear top-1 accuracy in Kinetics during pretraining for different forms of future prediction tasks: Vanilla, Guided, Masked, and different displacements. The longer the displacement the better the accuracy is. Note that this is an online evaluation where the model only sees 4 frames (that for ”Masked” and ”Guided” are distorted) hence performance is far below models evaluated on the whole 10s of video (250 frames), often with plenty of test time augmentation, but the relative performance between the various curves is informative.
@@ 5.3 IID vs Continuous Learning
Dummy baseline. For reference, we also include a dummy blind baseline, proposed in previous online learning works [16], which exploits the temporal correlations in the target stream without seeing the input frames at all: this ”model” outputs the mean of the previously seen pixels at the same locations for depth and pixel, and the most frequent colors for segmentation, for the previous target. It produces random performance for out-of-stream evaluation where consecutive targets are unrelated [^2] Another baseline that may be of interest is one that copies the previous input frames instead of target frames – this could do well for both train and eval for pixels, but poorly for segmentation and depth.

### v2

@@ Abstract
We introduce a framework for online learning from a single continuous video stream – the way people and animals learn, without mini-batches, data augmentation or shuffling. This poses great challenges given the high correlation between consecutive video frames and there is very little prior work on it. Our framework allows us to do a first deep dive into the topic and includes a collection of streams and tasks composed from two existing video datasets, plus methodology for performance evaluation that considers both adaptation and generalization. We employ pixel-to-pixel modelling as a practical and flexible way to switch between pre-training and single-stream evaluation as well as between arbitrary tasks, without ever requiring changes to models and always using the same pixel loss. Equipped with this framework we obtained large single-stream learning gains from pre-training with a novel family of future prediction tasks, found that momentum hurts, and that the pace of weight updates matters. The combination of these insights leads to matching the performance of IID learning with batch size 1, when using the same architecture and without costly replay buffers. An overview of the paper is available online at  https://sites.google.com/view/one-stream-video.
@@ 1 Introduction
Figure 1: Top: We introduce a framework for studying continuous learning in a single video stream. This is a natural yet unstudied problem, different from standard independent and identically distributed (IID) learning in video where batches contain clips from random videos in a random order. Bottom: We propose pixel-to-pixel models to evaluate our approach across prediction tasks (prediction of future frames, depth, segmentation). We measure both adaptation to the video stream – the model here updates its weights (learns) continuously to improve prediction – as well as generalization to out-of-stream clips – the model being adapted on the first stream is now evaluated on a different held-out stream without being allowed to adapt to it. We propose to maximize both adaptation and generalization.
@@
But what happens if we attempt to learn continuously from a single video stream (Figure 1), meaning batch size 1 and high frame rate (e.g. 25 fps)? We do not see thousands of dogs per second, unlike the standard ImageNet setting, where images of even a same dog breed are quite uncorrelated at several scales – capturing a wide diversity of edges, textures, etc. Standard deep learning approaches lead to smooth optimization there, but it is very different from the setting of learning from a single video stream, where things happen slowly, at much longer timescales, with great correlation or even no change between frames for periods of time; see Figure 2. Do standard deep learning tools work in this setting? This is currently unknown territory, with very few attempts touching it at all [33] – and it is the focus of this paper.
We are interested in this problem because models should be able to adapt after deployment to their environment, hence receiving a single stream of information. This is needed for embodied intelligence (robotics), but also for any type of digital assistant that we would want to be able to adapt to a specific user’s needs.
@@ 2 Related work
Online learning from a single video stream. One of the few papers dealing with this problem [33] proposed a minimum-redundancy “replay” buffer [27] to deal with temporal correlations. Replay buffers are not the sexiest research avenue for continual learning [41] and increase computation proportionally to the size of the buffer used. Additionally, in [33] performance on the video streams itself was never evaluated, only on unrelated image classification tasks. Test-time training is the next closest problem, such as presented in TTTVS [45] (also related  [48]). TTVS’s motivation is to “improve inference quality" by leveraging self-supervised test-time optimization. Our motivation is to propose single-stream learning, which is reflected in much longer streams (24h in this paper, compared to seconds or few minutes). One challenge in single-stream learning is how to fully exploit available hardware parallelism in a batch size 1 setting [6, 23, 24].
Learning from a single video. Also relevant are experiments training ConvNets [21] or ViTs [8] on a single image [1] or a long video [43]. While outside of the streaming setting – using data shuffling, augmentation and large batches, these papers showed that learning can be quite successful in this setting (especially for the shallower layers), which was quite encouraging.
Continual learning. The continual learning literature (e.g. EWC  [19]) tends to focus on learning a sequence of tasks, the goal being to learn new tasks as fast as possible without forgetting previous ones, such as in class incremental learning (e.g. iCarl  [36]) and CLEAR [22]. Most relevant is online continual learning, where models learn from a stream of data, visited once. One challenge is model cheating by exploiting labels in the data  [15], or by learning spurious features [46] that have limited generalisation power. Most of the work in this area has been on collections of images, rather than video.
@@
Representation learning to the rescue. Recent continual learning papers reported that some of the challenges of the problem are partially mitigated by using features pretrained in IID settings [26, 35, 44]. Various pretraining strategies exist: self-distillation (e.g. BYOL  [13], BRaVe [37], DINO [4]), contrastive (CLIP [34], VideoMoco [28], VITO [30]), and masked auto-encoding (e.g. VideoMAE [42], TubeViT [31], AudioVisualMAE [11], SiameseMAE [14],  [2]). We investigate this aspect in our setup.
@@ 3 The Framework
Overview. The following is repeated throughout a video stream: a (potentially pretrained) model processes $n$ frames (a time step) from an input video stream, and predicts frames for a future time step. Online (in-stream) performance metrics as well as an L2 loss function are then recorded by comparing predicted and target frames. Given the loss, gradients are computed and the model weights are updated by an optimizer. The target frames can come from the input stream itself or from an associated stream (e.g. for datasets having semantic segmentation or depth). In parallel with in-stream evaluation, we periodically assess performance on a held-out stream (composed of clips from the validation set of the dataset), to assess generalization.
3.1 Unified pixel-to-pixel modelling
@@
Replay buffer. We explored replay buffers [27], which keep a large cache of previous examples, then form batches by sampling from it. They have the disadvantage of increasing computational cost significantly over operating with batch size 1 (increases the theoretical computational cost of learning $K$ times, where $K$ is the batch size).
@@ 3.3 Evaluation
Learning a video stream task with a continuously updated model can be seen to decompose into: 1) discovering strong features that are general, for example learning that certain configuration of edges correspond to chair legs for semantic segmentation (generalization), which is likely to happen over long timescales. 2) by specializing to the particular scene being seen in the video, e.g. learning that the chair is the only red object in the scene and using redness to label the chair pixels (adaptation), which can happen over short timescales. Adaptation can be quite useful and lead to efficient visual mechanisms, but may not necessarily lead to generalization. Dummy models that only e.g. exploit temporal correlations in the target stream could misleadingly appear to perform very well [15]; see Table 4, Blind model.
We propose to evaluate both aspects by computing a score continuously in-stream, on the video stream task the model is learning from, to measure adaptation. Separately and periodically we compute the same score out-of-stream, on a held-out stream with the optimizer disabled (out-of-stream). We use the training sets of Ego4D and ScanNet to do in-stream evaluation, and the validation sets for out-of-stream evaluation.
Cumulative scores. Comparing single-stream learning models and approaches can be difficult if done naively. On the training stream, difficulty of the task may vary over time, for example certain parts of a stream may be more unpredictable due to camera motion or other factors, leading to natural oscillations. On the generalization side we care not just about performance at a particular moment in time, but also how fast it ramps up as training on the sequential video stream progresses. With this in mind, we use a global score over the whole stream – we do this by averaging the performance over 10,000 evenly-spaced points interpolated from the steps in both in-stream and out-of-stream settings.
@@ 4 Generalized Future Prediction
One of the most promising directions for continual learning, as identified in prior work, is representation learning. We are particularly interested in pixel-to-pixel approaches keeping consistent with the overall proposed framework where we never change neither the architecture nor the loss. We propose a family of pretraining methods that generalize future video prediction [40], and that follow a same pattern: given one input video clip the model is trained to predict a future clip from the same video (both clips are 4 frames long). We consider three variants, illustrated in Fig. 3:
  • The easiest one, Guided future prediction, replaces a few patches from the input clip with patches in the same position from the future clip, hence narrowing down the range of possibilities.
@@
  • Masked future prediction is a variation of Masked Auto-Encoding where the model must predict the future clip based on a partial view of the current clip.
Guided future prediction is related to Siamese MAE [14, 2, 47], but simpler, requiring a single forward pass over one model. Masked future prediction is related to video MAE-type approaches [42, 10] but with a strict separation between disjoint input and output clips, whereas in traditional video MAE there is a single clip that gets uniformly corrupted as part of denoising auto-encoding.
@@ 5.2 Pretraining
Figure 8: Linear top-1 accuracy in Kinetics during pretraining for different forms of future prediction tasks: Vanilla, Guided, Masked, and different displacements. The longer the displacement the better the accuracy is. Note that this is an online evaluation where the model only sees 4 frames (that for "Masked" and "Guided" are distorted) hence performance is far below models evaluated on the whole 10s of video (250 frames), often with plenty of test time augmentation, but the relative performance between the various curves is informative.
@@ 5.3 IID vs Continuous Learning
Dummy baseline. For reference, we also include a dummy blind baseline, proposed in previous online learning works [15], which exploits the temporal correlations in the target stream without seeing the input frames at all: this "model" outputs the mean of the previously seen pixels at the same locations for depth and pixel, and the most frequent colors for segmentation, for the previous target. It produces random performance for out-of-stream evaluation where consecutive targets are unrelated [^2] Another baseline that may be of interest is one that copies the previous input frames instead of target frames – this could do well for both train and eval for pixels, but poorly for segmentation and depth.